OpenStudy (anonymous):

1. Solve S = 4v2 for v

OpenStudy (lgbasallote):

first divide both sides by 4 \[\implies \frac S4 = v^2\] now..take the square root of both sides to isolate v

OpenStudy (anonymous):

How do I get rid of the square root?

OpenStudy (lgbasallote):

you take the square root..not get rid of it

OpenStudy (anonymous):

Well yeah. How do I do that?

OpenStudy (lgbasallote):

\[\sqrt{\frac S4} = \sqrt{v^2}\] like that

OpenStudy (lgbasallote):

then \[\sqrt{v^2} \implies v\] so \[v = \sqrt{\frac S4}\]

OpenStudy (anonymous):

so \[ v=s/4\]

OpenStudy (lgbasallote):

no. \[\sqrt{\frac S4}\]

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

m=2x + 3y for y

OpenStudy (anonymous):

I divide 3 right?